OpenStudy (anonymous):

If a snowball melts so that its surface area decreases at a rate of 0.2cm^2/min, find the rate at which the diameter decreases when the diameter is 9cm.

OpenStudy (bahrom7893):

Surface area of a sphere is 4pi*r^2 i believe

OpenStudy (bahrom7893):

Okay so: S = 4*pi*(d/2)^2 = pi*d dS/dt = pi*dd/dt

OpenStudy (bahrom7893):

wait typo

OpenStudy (bahrom7893):

S = pi*d^2 dS/dt = 2pi*d*dd/dt -0.2 = 2pi*9*dd/dt

OpenStudy (bahrom7893):

-0.2 = 18pi*(dd/dt) dd/dt = -0.2/(18pi)

OpenStudy (anonymous):

\[V=\frac{4}{3}\pi r^3 \implies S.A.=\frac{dV}{dr}=4\pi r^2\] The radius is half the diameter: \[\S.A.=\pi D^2 \implies \frac{d (S.A)}{dt}=2 \pi D \frac{d D}{dt} \implies \frac{d D}{dt}=\frac{1}{2\pi D}\frac{d (S.A)}{dt}\] Just plug in.